OpenStudy (anonymous):

Find the number c that satisfies the conclusion of the Mean Value Theorem. f(x)=x/x+2 Find the number c that satisfies the conclusion of the Mean Value Theorem. f(x)=x/x+2 @Mathematics

OpenStudy (anonymous):

guess there is something missing here. ...

OpenStudy (anonymous):

oh right sorry. the interval is [1,4]

OpenStudy (anonymous):

then you can do it. find \[f(4),f(1),\frac{f(4)-f(1)}{3}\] and also find \[f'(x)\] then set \[f'(x)=\frac{f(4)-f(1)}{3}\] and solve for x

OpenStudy (anonymous):

need help with the steps?

OpenStudy (anonymous):

\[f(4)=\frac{4}{4+2}=\frac{2}{3}\] \[f(1)=\frac{1}{1+2}=\frac{1}{3}\] \[f(4)-f(1)=\frac{1}{3}\] \[\frac{f(4)-f(1)}{3}=\frac{1}{9}\] check my arithmetic because i am typing it not write it

OpenStudy (anonymous):

also you need \[f'(x)=\frac{2}{(x+2)^2}\] and so now you set \[\frac{2}{(x+2)^2}=\frac{1}{9}\] and solve for x

OpenStudy (anonymous):

that's actually where my main problem is. i don't know how to solve for x at this point

OpenStudy (anonymous):

ooooooooooooooooh

OpenStudy (mathteacher1729):

You're looking for the value of x where the slope of your function is equal to the slope of the dotted line. :)

OpenStudy (anonymous):

solve as a ratio \[\frac{2}{(x+2)^2}=\frac{1}{9}\] \[(x+2)^2=18\]

OpenStudy (anonymous):

i know it should be 18=(x+2)^2

OpenStudy (anonymous):

yeah exactly

OpenStudy (anonymous):

then done in two simple steps \[x+2=\pm\sqrt{18}\] \[x=-2\pm3\sqrt{2}\] but only one is in your interval

OpenStudy (anonymous):

sometimes all it takes is asking and you see what you need to do before you get the answer

OpenStudy (anonymous):

by the way these questions stink, because mvt says nothing about "find the c" it just says there is one

OpenStudy (anonymous):

okay. thanks so much!! yeah they really do. do you think you could help me with 1 other one?

OpenStudy (anonymous):

post sure why not

OpenStudy (anonymous):

If f(3) = 7 and f '(x) ≥ 3 for 3 ≤ x ≤ 7, how small can f(7) possibly be?

OpenStudy (anonymous):

ok this question looks weird but you can actually do it in your head

OpenStudy (anonymous):

the very smallest f' can be is 3. that is like saying that you must go at least 3 units in the y for every unit in the x.

OpenStudy (anonymous):

so if f(3) = 7, then what is the least f(7) can be? well 7 is 4 units from 3, so you must have increased at least 3*4 = 12 units, and 12 more than 7 is 19 the reason this is confusing is that the moron who wrote this question used 3 and 7 repeatedly

OpenStudy (anonymous):

now we use mvt and write it out explicitly

OpenStudy (anonymous):

we know there is a number c such that \[\frac{f(7)-f(3)}{7-3}=f'(c)\geq 3\] \[\frac{f(7)-7}{4}\geq 3\] \[f(7)-7\geq 12\] \[f(7)\geq 19\]

OpenStudy (anonymous):

the way you said to to do it in your head makes perfect sense. thank you!!

OpenStudy (anonymous):

it would have been much easier to understand if it said \[f(2) =5,f'(x)\geq10\] and then asked what is the least \[f(8)\] can be

OpenStudy (anonymous):

you would say "from 2 to 8 is 6 units, and so we must have increased at least 6*10 = 60, and 60 more than 5 is 65"

OpenStudy (anonymous):

i completely understand that. again, thanks so much.

OpenStudy (anonymous):

yw